A point charge Q is placed at the center of a hemisphere . Find the electric flux passing through flat surface of hemisphere?

 

Grade:12th pass

1 Answers

ananthupotti
7 years ago
According to gauss law, flux through a closed surface=q/(epsilon).
Here charge enclosed by hemisphere=q/2
therefore, flux =q/2(epsilon)
 
.